Moving walkways are often found in airports. One moving walkway carries passengers up an incline of 30°, A single passenger of mass 72 kg stands on the walkway. The speed of the walkway is 0.51ms−1. Show that the rate at which the walkway does work on the passenger is about 200W.


1
Expert's answer
2021-03-02T18:04:45-0500

The weight of the passenger is


"W = mg"

where "m = 72kg" is they mass, and "g = 9.8N\/kg" is the gravitational acceleration. The component of the weight that acts along the walkway is:


"W_a = W\sin \theta"

where "\theta = 30\degree".

By definition, the power is:


"P = Wv"

where "v = 0.51m\/s" is the speed of motion. Thus, obtain:


"P = mgv\sin\theta\\
P = 72\cdot 9.8\cdot 0.51\cdot \sin 30\degree \approx 180W"

Answer. 180 W.